Suppose someone e-mails you 10 different files you need to save one-by-one to your computer. Suppose that in total you have 6 different empty folders on your computer.

(i) How many ways can you save the files into the 6 folders if the temporal order in which you save the individual files does not matter.

(ii) How many ways can you save the files into the 6 folders if the temporal order in which you save the individual files matters.


1
Expert's answer
2021-12-13T14:40:10-0500

i)

"N=\begin{pmatrix}
 n+ m-1\\
 m
\end{pmatrix}"

where m is number of files,

n is number of folders

"N=\begin{pmatrix}
 6+ m-1\\
 m
\end{pmatrix}=\begin{pmatrix}
 5+ m\\
 m
\end{pmatrix}=\frac{(5+m)!}{5!m!}"


ii)

"N=n^m=6^m"